### v1

@@ 3.3.2 Series Hybrid architecture
The Series Hybrid architecture (Fig. 1(b)) involves placing both the dry- and icy-regolith processing plants within the PSR to further process dry regolith after water extraction to obtain oxygen. Water extracted from icy regolith and generated through carbothermal reduction is transported to a PEL for electrolysis, liquefaction and storage, requiring the same water tanker system as the Parallel Hybrid architecture.
@@ 4 Lunar ISRU hybrid plant modeling
Fig. 2 shows the system architectures selected for this study. For each subsystem, a technology is modeled using ISRULib. ISRULib is an open-source component- and system-level library of ISRU models developed by the Technical University of Munich to carry out high-level technological trade-offs and preliminary architectural definitions [49]. To estimate the mass and power of the hybrid architectures proposed in this paper, parametric sizing models for carbothermal reduction and water extraction subsystems are newly developed and added to ISRULib. The models estimate mass and power budgets, as well as the performance of different technologies, based on analytical parametric calculations, extrapolated experimental results, surrogate models from numerical simulations, and already existing ISRU hardware.
@@
Figure 2: System architecture of the hybrid ISRU production plants. Mass flows across subsystems are shown. The diagram blocks crossed out in red are not modeled.
@@ 4.1 Carbothermal reduction of dry regolith
The reaction of silica with methane has been extensively modeled by Balasubramaniam et al. [52]. The reaction
@@
occurs at 1250-2000 C [53]. To achieve such high temperatures, Balasubramaniam et al. [52] and Linne et al. [7] proposed the use of solar thermal energy to melt a concentrated area of regolith. These studies also suggested to use regolith as a thermal insulator to protect reactor walls from molten regolith. The use of concentrated solar thermal energy has been experimentally simulated using lasers to demonstrate the CR process of lunar regolith simulants in vacuum conditions [10].
Figure 3 depicts a high-level schematic of the parametric sizing model used for CR reactor mass and power estimations. From silica content in the regolith, conversion rate, and dimensions of a molten regolith zone, the reaction time of one molten zone can be calculated using the model developed by Balasubramaniam et al. [52]. The estimated reaction time is used to calculate the minimum number of molten zones required to meet the target production rate. The number of molten zones can affect the dimensions of the reactor design, which affects the total mass of the reactor (see Fig. 4 for the design of the reactor).The main material of the reactor and its minimum allowable wall thickness is set to Inconel and 3.0 mm, respectively.
@@
The total power required from the solar concentrator can be divided into four: power for heating and melting the regolith, for carrying out the reaction, and the conductive and radiative heat losses to the environment. The heating power can be derived from the heat capacity of the regolith, which is modeled in the literature  [54, 55]. The latent heat of lunar highland regolith is calculated as 478.6 kJ/kg by Schreiner et al. [56]. The enthalpy of the reactions described by Equations 1 and 4 are reportedly 75.6 kJ/mol and 689.8 kJ/mol, respectively [57]. The conductive heat loss can be calculated from the geometry of the molten zone and the reactor, and the thermal conductivity of lunar regolith. Cremers and Hsia [58] reported the thermal conductivity of Apollo 16 highland regolith as $1.0-1.5\times 10^{-3}\mathrm{W/(m\cdot K)}$ at around 400 K. Due to this extremely low thermal conductivity of lunar regolith, the conductive heat loss can be practically ignored. The radiative heat loss can be calculated from the molten zone geometry and its temperature and the environment temperature. Following the assumptions made in [57], we use 0.7 as the emissivity of molten regolith, and we assumed that there is an insulation cover on each molten zone with a small hole on it. Due to these insulation covers, the radiation loss can be limited to 6.25% of the case without the covers. With these assumptions, the heat loss from one molten zone can be calculated as about 26.6 W.
The mass of the methanation chamber is modeled based on the work of Schrenk [59]. In this model, the mass of a Sabatier chamber is scaled based on a mass of those demonstrated in literature such as [60]. Following Ref. [59], the specific mass used for the methanation subsystem is 17.9 kg per 1.0 kg/h of water produced.
@@ 4.2 Water extraction from icy regolith
Figure 5 depicts a schematic of water extractors’ mass and power estimation developed for this paper. From the extraction efficiency and the target oxygen and hydrogen production rate, the target water extraction rate can be calculated. From the relationship between the input power to an extractor, the average water ice content in excavated regolith, and the water extraction rate, mentioned above, the model can estimate the required power. Note that the relationship is fit to a quadratic equation from the data points from [22], and the maximum allowable power input per reactor is set to 2500 W to avoid extrapolation on the higher power side. When the input power requirement exceeds 2500 W, the number of reactors is modified iteratively.
@@ 4.3.1 Excavation and handling subsystem
The excavator models are based on NASA’s Regolith Advanced Surface Systems Operations Robot (RASSOR) 2.0 [62]. RASSOR 2.0 includes two sets of two bucket drums that produce opposing digging forces, causing a net-zero horizontal reaction force and enabling its operation under reduced lunar gravity. To adapt the models to the conditions of each excavation site, the regolith density is modified according to depth and water ice content.
@@ 4.3.2 Beneficiation subsystem
A particle size separator is included to remove the coarse regolith fraction. Although the particle size effect on the carbothermal reduction of lunar regolith is not well-researched, Samouhos et al. [51] claim that the smaller particle size leads to higher efficiency in the reduction by increasing the relative surface area between regolith and carbon. Haas and Khalafalla [63] also reported a decreasing efficiency of carbothermal reduction around the silica particle size of 0.13 mm. Without further information on particle size effects on carbothermal reduction of lunar regolith, this paper assumes that the particle size separator removes particles with a diameter larger than 0.15 mm. The particle size separator is based on the model by Linne et al. [7], where a trough covered by a size-sorting grate dumps regolith from a hopper to an auger conveyor that transports the granular material to the next subsystem.
@@ 4.3.3 Water transportation
The model developed by Kleinhenz and Paz [23] is considered for the water tankers discussed in Section 3.3. The tankers are composed of a mobility platform and a payload, which includes a water tank, a battery, and a communication and navigation unit. The mobility platform mass is based on a payload ratio (payload mass/mobility platform mass) of 1.5, similar to RASSOR 2.0. The aluminum tank is modeled based on its water capacity, including a 50% ullage. The battery energy density is 140 Wh/kg, allowing for a maximum 80% discharge. These same battery properties are used for the RASSOR 2.0 excavators. The charging stations at the propellant production and H2O extraction plants can provide enough power to allow for a 5-hour excavator and a 10-hour tanker recharge time.
The tankers deliver water to a fixed aluminum tank at the H2/O2 production plant, which is 3 times larger than the total tankers’ water tank volume to account for operational mismatches. The water tank mass also includes a Commercial-Off-The-Shelf (COTS) pump , which is modeled based on those for space applications by Micropump, Inc. [26, 64]
@@ 4.4 Shared infrastructure
Most of the shared infrastructure between the oxygen and hydrogen coproduction from lunar icy and dry regolith is shown in green in Fig. 2. This shared infrastructure includes the water tank and electrolyzer, the H2 and O2 liquefaction and cryostorage subsystems, and high-pressure tanks to supply gas for the carbothermal reduction, accounting for buffer storage due to operational mismatches. A COTS pump is included before the pressurized vessels to store H2 and CH4 at 300 bar, and 200 bar, respectively [65], and average PEL temperature. It must be noted that the energy required to cool down the exhaust gas stream before the electrolysis is partially recuperated and used to preheat the intake methane stream. Therefore, an additional Condensing Heat Exchanger (CHX) is included between the pressurized vessel and the reactor to further increase the gas temperature to the operating temperature.
@@ 4.4.1 Electrolysis subsystem
Table 2 lists the mass and power consumption estimated for both systems sized to generate 10 t of oxygen and 1.25 t of hydrogen per lunar year. The mass and power in this table are summations of all components shown in Fig. 6. As can be seen in Table 2, the mass of the PEM electrolysis subsystem can be significantly lighter, whereas the SOXE electrolysis requires less power. As further discussed in Section 5.1, this difference in mass between these two systems can be trivial compared to the entire system mass. Conversely, the electrolysis subsystem is one of the main power drivers, and the difference between these two technologies (i.e., $\sim$16 kW) can have a significant impact. Therefore, this paper employed the SOXE subsystem for further analysis.
@@
Mass (kg) Power (kW)
PEM 9.1 23.00
SOXE 100.6 7.18
@@ 5.1 Comparison of Resource Extraction Architecture
Element CR WE Parallel∗2 Series CR WE Parallel∗2 Series
@@
   Other dry regolith processing subsystems∗3 33 - 32 29 2 - 1 1
@@
H2/ O2 production and storage
   SOXE electrolysis subsystem 101 101 101 101 7,183 7,183 7,183 7,183
   Liquefaction subsystem 192 1,870 1,039 680 5,616 28,093 16,855 12,038
   Storage and recycle subsystems 348 551 518 450 71 0 69 70
@@
   Additional H2 and tankssubsystems∗5 2,161 - 1,101 1,664 - - - -
   Additional CH4 and tankssubsystems∗6 966 - 491 746 - - - -
Total 4,265 5,872 5,051 6,988 79,515 63,596 72,010 77,892
$\ast 1$: All values include 30% growth [45] except for additional H2 mass.
@@
$\ast 3$: Supply and discharge hoppers, and particle size separators.
$\ast 4$: Supply and discharge hoppers.
$\ast 5$: Additional H2 for reactant loss (91% recycling efficiency) and product compensation.
$\ast 6$: 91% recycling efficiency.
Figures 7 and 8 are bar charts of mass and power breakdown. In addition to the four architectures in Table 3, the Parallel Hybrid architectures with different hybrid ratios are added. The landed mass of the CR architecture turned out to be the lightest compared to the others for one year of operation (Fig. 7(a)). As the operational period becomes longer, the additional hydrogen and methane from Earth become heavier, making the WE architecture the lightest (Fig. 7(b)). The main mass drivers other than additional methane and hydrogen are the water tankers and the liquefaction subsystem. Since liquefying hydrogen requires heavy cryocoolers (see, e.g., Deserranno et al. [73]), the liquefaction subsystem of the CR architecture becomes the lightest. The Series Hybrid architecture turns out to be the heaviest in both analyzed operational periods. This is because the Series Hybrid architecture largely depends on oxygen from CR rather than WE, requiring additional hydrogen and methane from Earth. Furthermore, since all water is produced in the PSR, this architecture requires the same amount of water tankers as the WE dedicated plant, which is another major mass driver.
@@
Figure 8 is the bar chart of the power breakdown. The main power drivers are the CR reactor, the water extractor, the electrolysis subsystem, and the liquefaction subsystem. As can be seen, the water extractor itself requires significantly less power than the CR reactor. However, due to the liquefaction of hydrogen, the total difference in power consumption between the CR and WE architectures is about 16 kW.

### v2

@@ 3.3.2 Series hybrid architecture
The SH architecture (Fig. 1(d)) involves placing both the dry- and icy-regolith processing plants within the PSR to further process dry regolith after water extraction to obtain oxygen. Water extracted from icy regolith and generated through carbothermal reduction is transported to a PEL for electrolysis, liquefaction and storage, requiring the same water tanker system as the PH architecture.
@@ 4 Lunar ISRU hybrid plant modeling
Fig. 3 shows the flow diagrams of the system architectures studied in this paper. For several subsystems, technologies are modeled using ISRULib. ISRULib is an open-source component- and system-level library of ISRU models developed by the Technical University of Munich to carry out high-level technological trade-offs and preliminary architectural definitions [51]. To estimate the mass and power of the hybrid architectures proposed in this paper, parametric sizing models for carbothermal reduction and water extraction subsystems are newly developed. The models estimate mass and power budgets, as well as the performance of different technologies, based on analytical parametric calculations, surrogate models from numerical simulations, and already existing ISRU hardware. Table 2 lists subsystem models employed in this paper. Each of them is explained below in detail.
(a) Carbothermal reduction architecture.
(b) Direct water extraction architecture.
Figure 2: Flow diagrams of ISRU architectures considered in this paper. Mass flows across subsystems are shown. The diagram blocks crossed out in red are not modeled.
@@
Figure 3: (Continued.)
Table 2: Summary of subsystem model types and assumptions
Element Model type Notes and assumptions Refs.
Regolith mining
   Excavator Existing hardware RASSOR 2.0 [52]
Dry regolith processing
   CR reactor Param. size modeling Material: Inconel [53, 54, 55]
   Methanation reactor Specific mass Reference: produced H2O, kg/h [56]
   Particle size separator Param. size modeling Material: aluminum [57]
   Condensor heat exchanger Param. size modeling Material: titanium [58]
   Hopper Param. size modeling Material: Inconel (discharge and disposal)
Material: aluminum (supply)
Icy regolith processing
   H2O extractor Surrogate modelling from (Kiewiet et al.) Material: Inconel [22]
and param. size modeling
   H2O tank Param. size modeling Material: aluminum
   Hopper Param. size modeling Material: aluminum
Mobile H2O transportation
   H2O tanker Param. size modeling Based on payload ratio, aluminum tank,
and energy density
Electrolysis
   PEM electrolysis stack Param. size modeling [56]
   SOXE electrolysis stack Param. size modeling [59]
   Dryers Param. size modeling Material: titanium [56]
   Heat exchanger Param. size modeling Material: titanium [58]
   Pump Specific mass Reference: fluid flow rate, L/min [60]
H2/ O2 liquefaction
   H2 Liquefaction subsystem Existing hardware, Compressor (Deserranno, et al.) [61]
and param. size modeling material: aluminum and neon [58]
   O2 Liquefaction subsystem Existing hardware, Compressor (CryoTel®DS 30) [62]
and param. size modeling material: aluminum and neon [58]
Recycle and storage
   Gas storage tank Param. size modeling Material: carbon/epoxy
   Cryogenic storage tank Param. size modeling Material: aluminum and MLIs [63]
   Pump Specific mass Same as above -
Thermal management
   Radiators Specific mass Reference: reject heat, kg/kW [32, 64]
param. = parametric
@@ 4.1 Carbothermal reduction of dry regolith
The reaction of silica with methane has been extensively modeled by Balasubramaniam et al. [53] The reaction
@@
occurs at 1250-2000 C [67]. To achieve such high temperatures, the use of solar thermal energy to melt a concentrated area of regolith has been proposed and demonstrated [68, 69]. These studies also suggested to use regolith as a thermal insulator to protect reactor walls from molten regolith. The use of concentrated solar thermal energy has been experimentally simulated using lasers to demonstrate the CR process of lunar regolith simulants in vacuum conditions as well [10].
Figure 4 depicts a high-level schematic of the parametric sizing model used for CR reactor mass and power estimations. From silica content in the regolith, conversion rate, and dimensions of a molten regolith zone, the reaction time of one molten zone can be calculated using the model developed by Balasubramaniam et al. [53] The estimated reaction time is used to calculate the minimum number of molten zones required to meet the target production rate. The number of molten zones can affect the dimensions of the reactor design, which affects the total mass of the reactor (see Fig. 5 for the design of the reactor).The main material of the reactor and its minimum allowable wall thickness is set to Inconel and 3.0 mm, respectively.
@@
The total power required from the solar concentrator can be divided into four: power for heating and melting the regolith, for carrying out the reaction, and the conductive and radiative heat losses to the environment. The heating power can be derived from the heat capacity of the regolith, which is modeled in the literature  [54, 55]. The latent heat of lunar highland regolith is calculated as 478.6 kJ/kg by Schreiner et al. [70]. The enthalpy of the reactions described by Equations 1 and 4 are reportedly 75.6 kJ/mol and 689.8 kJ/mol, respectively [71]. The conductive heat loss can be calculated from the geometry of the molten zone and the reactor, and the thermal conductivity of lunar regolith. Cremers and Hsia [72] reported the thermal conductivity of Apollo 16 highland regolith as $1.0-1.5\times 10^{-3}\mathrm{W/(m\cdot K)}$ at around 400 K. Due to this extremely low thermal conductivity of lunar regolith, the conductive heat loss can be practically ignored. The radiative heat loss can be calculated from the molten zone geometry and its temperature and the environment temperature. Following the assumptions made in [71], we use 0.7 as the emissivity of molten regolith, and we assumed that there is an insulation cover on each molten zone with a small hole on it. Due to these insulation covers, the radiation loss can be limited to 6.25% of the case without the covers. With these assumptions, the heat loss from one molten zone can be calculated as about 26.6 W. From these parameters and batch processing time, the average total power required for CR reactors is calculated.
The mass of the methanation chamber is modeled based on the work of Schrenk [56]. In this model, the mass of a Sabatier chamber is scaled based on a mass of those demonstrated in literature such as [73]. Following [56], the specific mass used for the methanation subsystem is 17.9 kg per 1.0 kg/h of water produced. As discussed later in Section 5.1, the mass of the methanation chambers is not a main mass driver, and therefore, this simple specific mass model is enough considering the scope of this paper. Parametric sizing models for different chamber designs would be desirable for more detailed CR architecture design decision-making.
@@ 4.2 Water extraction from icy regolith
Figure 6 depicts a schematic of water extractors’ mass and power estimation developed for this paper. From the extraction efficiency and the target oxygen and hydrogen production rate, the target water extraction rate can be calculated. From the relationship between the input power to an extractor, the average water ice content in excavated regolith, and the water extraction rate, mentioned above, the model can estimate the required power. Note that the relationship is fit to a quadratic equation from the data points from Kiewiet et al. [22], and the maximum allowable power input per reactor is set to 2500 W to avoid extrapolation on the higher power side. When the input power requirement exceeds 2500 W, the number of reactors is modified iteratively.
@@ 4.3.1 Excavation and handling subsystem
The excavator models are based on NASA’s Regolith Advanced Surface Systems Operations Robot (RASSOR) 2.0 [52]. RASSOR 2.0 includes two sets of two bucket drums that produce opposing digging forces, causing a net-zero horizontal reaction force and enabling its operation under reduced lunar gravity. It should be noted that excavation of icy regolith may be more challenging than that of dry regolith depending on how ice is formed in a PSR [75]. Without having enough information about the water ice structure on the Moon, this paper assumes RASSOR 2.0 is used for both dry and icy regolith excavation. The required power for the excavation is also assumed to be the same between dry and icy regolith. While the first generation of RASSOR demonstrated the excavation of lunar regolith simulant with water ice [76], further demonstration for various water ice content needs to be conducted.
@@ 4.3.2 Beneficiation subsystem
A particle size separator is included to remove the coarse regolith fraction. Although the particle size effect on the carbothermal reduction of lunar regolith is not well-researched, Samouhos et al. [66] claim that the smaller particle size leads to higher efficiency in the reduction by increasing the relative surface area between regolith and carbon. Haas and Khalafalla [77] also reported a decreasing efficiency of carbothermal reduction around the silica particle size of 0.13 mm. Without further information on particle size effects on carbothermal reduction of lunar regolith, this paper assumes that the particle size separator removes particles with a diameter larger than 0.15 mm. The particle size separator is based on the model by Linne et al. [7], where a trough covered by a size-sorting grate dumps regolith from a hopper to an auger conveyor that transports the granular material to the next subsystem. The main mass and power driver of this beneficiation subsystem is the auger, which is modeled based on Benedi et al. [57]. Hence, while the size sorter in the study by Linne et al. is designed for removing particles larger than 25 mm, the differences in the mass and power consumption between their system and the proposed one are assumed trivial. Note that there may be more appropriate techniques for the size sorting of small highland regolith particles. For instance, Al Moinee et al. [78] used an electrostatic sieve to separate particles smaller than 0.5 mm into three groups. For more detailed ISRU design consideration, a tradeoff study of available beneficiation technologies and a careful consideration of feasible operations would be necessary.
@@ 4.3.3 Water transportation
The model developed by Kleinhenz and Paz [23] is considered for the water tankers discussed in Section 3.3. The tankers are composed of a mobility platform and a payload, which includes a water tank, a battery, and a communication and navigation unit. The mobility platform mass is based on a payload ratio (payload mass/mobility platform mass) of 1.5, similar to RASSOR 2.0. The aluminum tank is modeled based on its water capacity, including a 50% ullage. The battery energy density is 140 Wh/kg, allowing for a maximum 80% discharge. These same battery properties are used for the RASSOR 2.0 excavators. The charging stations at the propellant production and H2O extraction plants can provide enough power to allow for a 5-hour excavator and a 10-hour tanker recharge time. As mentioned in Sec. 3.3.1, water starts to thaw after a tanker leaves the PSR. The required time for this process is calculated from the thermal environment of the tanker. The developed model compares this time with the charge time and adds extra time for waiting for this process to be done when necessary.
The tankers deliver water to a fixed aluminum tank at the H2/O2 production plant, which is 3 times larger than the total tankers’ water tank volume to account for operational mismatches. The water tank mass also includes a Commercial-Off-The-Shelf (COTS) pump , which is modeled based on those for space applications by Micropump, Inc. [26, 58, 60]
@@ 4.4 Shared infrastructure
Most of the shared infrastructure between the oxygen and hydrogen coproduction from lunar icy and dry regolith is shown in green dashed lines in Fig. 3. This shared infrastructure includes the water tank and electrolyzer, the H2 and O2 liquefaction and cryostorage subsystems, and high-pressure tanks to supply gas for the carbothermal reduction, accounting for buffer storage due to operational mismatches. A COTS pump is included before the pressurized vessels to store H2 and CH4 at 300 bar, and 200 bar, respectively [79], and average PEL temperature. It must be noted that the energy required to cool down the exhaust gas stream before the electrolysis is partially recuperated and used to preheat the intake methane stream. Therefore, an additional Condensing Heat Exchanger (CHX) is included between the pressurized vessel and the reactor to further increase the gas temperature to the operating temperature. Optimal storage conditions of reactants should be discussed in the future.
@@
4.4.1 Thermal management subsystem
Some subsystems require thermal management to exhaust heat. In this paper, radiators are considered to accomplish this purpose. Subsystems that require radiators in this paper are a CHX for CR, an electrolysis subsystem, and a liquefaction subsystem.
Due to different thermal conditions between a PSR and a PEL, the required surface area to reject a unit amount of heat should be different. Guerrero-Gonzalez and Zabel [32] estimated 40 kg of radiator required to reject 1 kW heat based on Guerrero-Gonzalez [81] at a PSR. Simonsen et al. [64] also estimated the required mass of a radiator system for different regions on the Moon. In their study, to reject 201 kW heat, the mass of the thermal management system was estimated as 10, 681 kg at the South Pole. Therefore, in this paper, the mass factor of 53 kg/kW is considered for radiators in a PEL.
Since CR is conducted in a PSR for the SH architecture, a radiator is assumed to be placed in a PSR. The rest of the radiators for the SH architecture as well as those of the PH architecture are placed in a PEL.
@@ 4.4.2 Electrolysis subsystem
Table 3 lists the mass and power consumption estimated for both systems sized to generate 10 t of oxygen and 1.25 t of hydrogen per lunar year. The mass and power in this table are summations of all components shown in Fig. 7 and radiators. As can be seen in Table 3, the mass of the PEM electrolysis subsystem can be significantly lighter, while it requires larger radiators. Regarding required power, the SOXE electrolysis requires less power. Therefore, this paper employed the SOXE subsystem for further analysis.
@@
SOXE PEM
Mass [kg] 418 567
   Electrolysis 101 12
   Termal management 317 555
Power [W] 7,756 23,000
@@ 5.1 Comparison of resource extraction architecture
Element CR WE PH∗2 SH CR WE PH∗2 SH
@@
   Particle size separator 5 - 2 2 15 - 8 10
   Condensor heat exchanger 4 - 2 3 - - - -
   Hoppers 31 - 31 27 15 - 8 10
@@
Electrolysis
   SOXE electrolysis stacks 100 100 100 100 7,182 7,182 7,182 7,182
   Dryers 1 1 1 1 - - - -
   Heat exchangers 0.2 0.2 0.2 0.2 - - - -
   Pumps 0.4 0.4 0.4 0.4 574 574 574 574
H2/ O2 liquefaction
   H2 Liquefaction subsystem - 1,665 834 475 - 22,113 11,057 6318
   O2 Liquefaction subsystem 179 179 179 179 5,616 5,616 5,616 5,616
Recycle and storage
   CH4 storage 27 - 27 27 - - - -
   CH4 recycling pump 0.4 0 0.4 0.4 67 - 67 67
   H2 storage 39 757 431 231 - - - -
   H2 recycling pump 0.4 0 0.4 0.4 5 - 2 3
   O2 storage 804 804 804 804 - - - -
Thermal management
   Radiators at PEL 513 453 483 439 - - - -
   Radiators at PSR - - - 45 - - - -
@@
   Additional H2∗3 1,591 - 795 1,220 - - - -
   Additional H2 tanks 571 - 306 445 - - - -
   Additional CH4∗4 906 - 453 696 - - - -
   Additional CH4 tanks 61 - 38 51 - - - -
Total 5,378 7,309 6,298 8,106 55,732 63,806 60,223 57,361
$\ast 1$: All values include 30% growth [47] except for additional H2 and CH4 mass.
@@
$\ast 3$: Additional H2 for reactant loss (91% recycling efficiency) and product compensation.
$\ast 4$: 91% recycling efficiency.
Figures 8 and 9 are bar charts of mass and power breakdown. In addition to the four architectures in Table 4, the Parallel Hybrid architectures with different hybrid ratios are added. The landed mass of the CR architecture turned out to be the lightest compared to the others for one year of operation (Fig. 8(a)). As the operational period becomes longer, the additional hydrogen and methane from Earth become heavier, making the WE architecture the lightest (Fig. 8(b)). The main mass drivers other than additional methane and hydrogen are the water tankers and the liquefaction subsystem. Since liquefying hydrogen requires heavy cryocoolers (see, e.g., Deserranno et al. [61]), the liquefaction subsystem of the CR architecture becomes the lightest. The SH architecture turns out to be the heaviest in both analyzed operational periods. This is because the SH architecture largely depends on oxygen from CR rather than WE, requiring additional hydrogen and methane from Earth. Furthermore, since all water is produced in the PSR, this architecture requires the same amount of water tankers as the WE dedicated plant, which is another major mass driver.
@@
As can be seen in Table 4 and Fig. 9, the main power drivers are the CR reactor, the water extractor, the electrolysis subsystem, and the liquefaction subsystem. As can be seen, the water extractor itself requires less power than the CR reactor. However, due to the liquefaction of hydrogen, the total difference in power consumption between the CR and WE architectures is about 8 kW. While the CR architecture requires the least power, the difference in power consumption between the CR and SH architectures is relatively small ($\sim 1.6$ kW). In the SH architecture, the low temperature at the PSR increases the energy required to heat regolith and leads to higher radiative heat loss. However, by partially relying on the less power-consuming WE process for oxygen production, the total power required for the CR reactor remains lower than that of the SH architecture.
It is important to note that the total power consumption is calculated under the assumption that water extraction and CR reactor operation occur simultaneously in hybrid architectures. If these processes are carried out at different times, power consumption at any given moment may be lower than the values presented here. Future studies should explore optimal operational strategies to minimize power consumption for each architecture.